### v1

# A replica analysis of under-bagging

Abstract
A sharp asymptotics of the under-bagging (UB) method, which is a popular ensemble learning method for training classifiers from an imbalanced data, is derived and used to compare with several other standard methods for learning from imbalanced data, in the scenario where a linear classifier is trained from a binary mixture data. The methods compared include the under-sampling (US) method, which trains a model using a single realization of the subsampled dataset, and the simple weighting (SW) method, which trains a model with a weighted loss on the entire data. It is shown that the performance of UB is improved by increasing the size of the majority class, even if the class imbalance can be large, especially when the size of the minority class is small. This is in contrast to US, whose performance does not change as the size of the majority class increases, and SW, whose performance decreases as the imbalance increases. These results are different from the case of the naive bagging in training generalized linear models without considering the structure of class imbalance, indicating the intrinsic difference between the ensembling and the direct regularization on the parameters.
1 Introduction
In the context of binary classification, class imbalance refers to the situations where one class overwhelms the other in terms of the number of data points. Learning on class imbalanced data often occurs in real-world classification tasks, which occur in diverse domains such as the medical data analysis (Cohen et al., 2006), the image classification tasks (Khan et al., 2018), and so on. The primary goal when training classifiers on such imbalanced data is to achieve good generalization to both minority and majority classes. However, standard supervised learning without any special care often results in poor generalization performance for the minority class. Therefore, special treatments are necessary and have been one of the major research topics in the machine learning (ML) community for decades.
Under-bagging (UB) (Wallace et al., 2011) is a popular and efficient method for dealing with a class imbalance that combines under-sampling (US) and bagging (Breiman, 1996). The basic idea of UB is to address label imbalance by randomly discarding a portion of the majority class data, ensuring an equal number of data points between the minority and majority classes. However, this data reduction increases the variance of the estimator since the total number of training data is reduced. To mitigate this, UB aggregates estimates for different under-sampled instances, similar to the aggregation performed in bootstrap aggregation (Breiman, 1996). US corresponds to a method of training a classifier on a single undersampled data. (Wallace et al., 2011) argues that UB is particularly effective in the training of overparametrized models, where the number of the model parameters far exceeds the size of the training data.
There are also another line of approach, that is, cost-sensitive methods, which modify the training loss to account for varying class penalties. However, it has been recently pointed out that training overparametrized models using cost-sensitive methods requires delicate design of loss functions (Lin et al., 2017; Khan et al., 2018; Byrd & Lipton, 2019; Ye et al., 2020; Kini et al., 2021) instead of the naive weighted class-entropy and so on. In contrast to such delicate cost-sensitive approaches, UB has the advantage of being relatively straightforward to use, as it achieves complete class balance in each under-sampled dataset.
Bagging is a natural approach to reduce the increased variance of weak learners, resulting from the smaller data size due to the resampling. However, unlike decision trees, which are the primary models used in bagging, neural networks can use simple regularization methods such as ridge regularization to reduce the variance. In fact, is has been experimentally reported that naive bagging, which ignores the structure of class balance, does not improve the classification performance of deep neural networks (Nixon et al., 2020). Furthermore, in generalized linear models (GLMs), it has been both theoretically (Sollich & Krogh, 1995; Krogh & Sollich, 1997; LeJeune et al., 2020; Du et al., 2023) and experimentally (Yao et al., 2021) reported that bagging with a specific sampling rate is exactly equivalent to the ridge regularization with properly adjusted regularization parameter. These results indicate that naive bagging plays the role of an implicit regularization in linear models. Given these results, the question arises whether it is better to use UB, which requires an increased computational cost proportional to the number of under-sampled datasets, when training linear models or neural networks, rather than just employing ridge regularization.
Taking into account the aforementioned situation, this study aims to clarify whether bagging or ridge regularization should be used when learning from class-imbalanced data. Specifically, we compare the performance of classifiers obtained by UB with those obtained by using a single under-sampled data with appropriate ridge regularization, which we refer to as the simple weighting (SW) method, when training a linear model on binary mixture data, through the $F$-measure. To do this, we derive a sharp characterization of the classifiers obtained by the randomly subsampled data in the asymptotic limit where the input dimension and the parameter dimension diverge at the same rate, and then use it to analyze the performance of UB, US and SW.
The contributions of this study are summarized as follows:
  • A sharp characterization of the statistical behavior of the linear classifier is derived, in which the classifier is obtained by minimizing the reweighted empirical risk function on binary mixture data, in the asymptotic limit where input dimension and data size diverge proportionally (Claim 4). This includes the properties of the estimator obtained from the under-sampled data and the simple weighted loss function. There, statistical property of the logits for a new input is described by a small finite set of scalar quantities determined as a solution of nonlinear equations, which we refer to as self-consistent equations (Claim 1). The sharp characterization of $F$-measure is also obtained (Claim 5). The derivation is based on the replica method of statistical mechanics (Mézard et al., 1987; Charbonneau et al., 2023; Montanari & Sen, 2024).
  • By using the derived sharp asymptotics, it is shown that the estimators obtained by UB have higher generalization performance in terms of $F$-measure than those obtained by US or SW. Although the $F$-measure is improved by increasing the size of the majority class while keeping the size of the minority class fixed in UB, the performance of US does not depend on the size of the excess examples of the majority class. These are clearly different from the case of the naive bagging in training GLMs without considering the structure of class imbalance. That is, the ensembling and the regularization are different when the data structure is taken into account. The performance of SW is even worse, i.e., its performance on the $F$-measure degrades as the size of the majority class increases, leading to a catastrophic performance degradation when training from data with small minority class.
  • Furthermore, it is observed that UB is robust to the existence of the interpolation phase transition where data from linearly separable to inseparable training data, although the performance of the standard interpolator obtained by a single realization of training data is affected by such a phase transition as reported in (Mignacco et al., 2020).
1.1 Further related works
Learning under class imbalance is a classic and notorious problem in data mining and ML. In fact, the development of dealing with imbalanced data was selected as “10 CHALLENGING PROBLEMS IN DATA MINING RESEARCH” (YANG & WU, 2006) about 20 years ago in data mining community. Due to its practical relevance, plethora of methodological research has been continuously conducted and are summarized in reviews (He & Garcia, 2009; Longadge & Dongre, 2013; Johnson & Khoshgoftaar, 2019).
Learning methods for imbalanced data can broadly be classified into two categories: cost- and data-level ones. In cost-sensitive methods, the cost function itself is modified to assign different costs to each class. A classical method includes weighting method that assigns a different constant multiplier for each data point depending on the class it belongs to. However, in the training of overparametrized models, it has been pointed out that a classical weighting methods fails and a more delicate handling of loss design is required to properly adjust the margins between classes (Lin et al., 2017; Khan et al., 2018; Byrd & Lipton, 2019; Ye et al., 2020; Kini et al., 2021). In data-level methods, the training data itself is modified to create a class-balanced dataset using resampling methods. The major approaches include US, which randomly discard a portion of the majority class data, and SMOTE (Chawla et al., 2002), which creates synthetic minority data points. Although the literature (Wallace et al., 2011) states that a combination of US and bagging is effective in learning overparametrized models, the comparison of other regularization techniques are not discussed.
From a technical viewpoint, our work is an analysis of the sharp asymptotics of resampling methods in which the input dimension and the size of the dataset diverge at the same rate. Analyzing such a sharp asymptotics of ML methods in this asymptotic regime has been actively conducted since the 1990s, using techniques from statistical mechanics (Opper & Kinzel, 1996; Opper, 2001; Engel & Van den Broeck, 2001), and it has recently become one of the standard analytical approaches. The technical tools for analyzing such asymptotics include the replica method (Mézard et al., 1987; Gerace et al., 2020; Charbonneau et al., 2023; Montanari & Sen, 2024), convex Gaussian minimax theorem (Thrampoulidis et al., 2018), random matrix theory (Mei & Montanari, 2022), adaptive interpolation method (Barbier & Macris, 2019; Barbier et al., 2019), approximate message passing (Sur & Candès, 2019; Feng et al., 2022), and so on.
Resampling methods have also been studied in the same proportional asymptotics. (Sollich & Krogh, 1995; Krogh & Sollich, 1997; Malzahn & Opper, 2001; 2002; 2003; LeJeune et al., 2020; Ando & Komaki, 2023; Bellec & Koriyama, 2024; Clarté et al., 2024; Patil & LeJeune, 2024) analyzes the performance of ensemble average of the linear models, and (Takahashi, 2023) analyzes the distribution of the ensemble average of the de-biased elastic-net estimator (Javanmard & Montanari, 2014b; a). However, these analyses do not address the structure of class imbalance. Ensembling in random feature models as a method to reduce the effect of random initialization are also considered in (D’Ascoli et al., 2020; Loureiro et al., 2022). Furthermore, the literature (Obuchi & Kabashima, 2019; Takahashi & Kabashima, 2019; 2020) derives the properties of the estimators trained on the resampled dataset in the analyses of the dynamics of approximate message passing algorithms that yields exact result in Gaussian or rotation-invariant feature data.
Notation Description
$N$ input dimension
$M^{+}$ size of the positive examples (assumed to be the minority class)
$M^{-}$ size of the negative examples (assumed to be the majority class)
$M$ $M^{+}+M^{-}$, the size of the overall training data
$\mu,\nu$ indices of data points
$i,j$ indices of the weight vector $\mathbf{\boldsymbol{w}}$
$[n]$ for an positive integer $n$, the set $\{1,\dots,n\}$
$v_{i}$ $i$-th element of a vector $\mathbf{\boldsymbol{v}}$
$\mathbf{\boldsymbol{x}}\cdot\mathbf{\boldsymbol{y}}$ for vectors $\mathbf{\boldsymbol{x}},\mathbf{\boldsymbol{y}}\in\mathbb{R}^{N}$, the inner product of them: $\mathbf{\boldsymbol{x}}\cdot\mathbf{\boldsymbol{v}}=\sum_{i=1}^{N}x_{i}y_{i}$.
$\mathbf{\boldsymbol{1}}_{N}$ an $N$-dimensional vector $(1,1,\dots,1)\in\mathbb{R}^{N}$
$\mbox{\rm 1}\mbox{\rm l}(\cdot)$ indicator function
${\mathcal{N}}(\mu,\sigma^{2})$ Gaussian density with mean $\mu$ and variance $\sigma^{2}$
${\rm Poisson}(\mu_{\ast})$ Poisson distribution with mean $\mu_{\ast}$
iid independent and identically distributed
$\mathbb{E}_{X\sim p_{X}}[f(X)]$ Expectation regarding random variable $X$
where $p$ is the density function for the random variable $X$
(lower subscript $X\sim p_{X}$ or $p_{X}$ can be omitted if there is no risk of confusion)
${\mathbb{V}}_{X\sim p_{X}}[f(X)]$ Variance: $\mathbb{E}[f(X)^{2}]-\mathbb{E}[f(X)]^{2}$
(lower subscript $X\sim p_{X}$ or $p_{X}$ can be omitted if there is no risk of confusion)
$\delta_{\rm d}(\cdot)$ Dirac’s delta function
Table 1: Notations
1.2 Notations
Throughout the paper, we use some shorthand notations for convenience. We summarize them in Table 1.
2 Problem setup
This section presents the problem setup and our interest in this work. The assumptions on the data generation process are first described, and then the estimator of the interest is introduced.
Let $D^{+}=\{(\mathbf{\boldsymbol{x}}_{\mu}^{+},1)\}_{\mu=1}^{M^{+}},\mathbf{\boldsymbol{x}}_{\mu}^{+}\in\mathbb{R}^{N}$ and $D^{-}=\{(\mathbf{\boldsymbol{x}}_{\nu}^{-},-1)\}_{\nu=1}^{M^{-}},\mathbf{\boldsymbol{x}}_{\nu}^{-}\in\mathbb{R}^{N}$ be the sets of positive and negative examples; in total we have $M=M^{+}+M^{-}$ examples. In this study, the positive class is assumed to be the minority class, that is, $M^{+}<M^{-}$. Also, it is assumed that positive and negative samples are drawn from a mixture models whose centroids are located at ±$\pm\mathbf{\boldsymbol{v}}/\sqrt{N}$ with $\mathbf{\boldsymbol{v}}\in\mathbb{R}^{N}$ being a fixed vector as follows:
$\mathbf{\boldsymbol{x}}_{\mu}^{+}$ $=\frac{1}{\sqrt{N}}\mathbf{\boldsymbol{v}}+\mathbf{\boldsymbol{z}}_{\mu}^{+},$ (1)
$\mathbf{\boldsymbol{x}}_{\nu}^{-}$ $=-\frac{1}{\sqrt{N}}\mathbf{\boldsymbol{v}}+\mathbf{\boldsymbol{z}}_{\nu}^{-},$ (2)
where $\mathbf{\boldsymbol{z}}_{\mu}^{+}$ and $\mathbf{\boldsymbol{z}}_{\nu}^{-}$ are noise vectors. We assume that $z_{\mu,i}^{+},z_{\nu,i}^{-}\sim_{\rm iid}p_{z},\mu\in[M^{+}],\nu\in[M^{-}],i\in[N]$, where the mean and the variance of $p_{z}$ are zero and $\Delta$, respectively, and also the higher order moments are finite. From the rotational symmetry, we can fix the direction of the vector $\mathbf{\boldsymbol{v}}$ as $\mathbf{\boldsymbol{v}}=(1,1,\dots,1)^{\top}$ without loss of generality. The goal is to obtain a classifier that generalize well to both positive and negative inputs from $D=D^{+}\cup D^{-}$. We focus on the training of the linear model, i.e., the model’s output $f(\mathbf{\boldsymbol{x}})$ at an input $\mathbf{\boldsymbol{x}}$ is a function of a linear combination of the weight $\mathbf{\boldsymbol{w}}$ and the bias $B$:
$f_{\rm model}(\mathbf{\boldsymbol{x}})=f\left(\frac{1}{\sqrt{N}}\mathbf{\boldsymbol{x}}\cdot\mathbf{\boldsymbol{w}}+B\right),$ (3)
where the factor $1/\sqrt{N}$ is introduced to ensure that the logit $\mathbf{\boldsymbol{x}}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+B$ should be of ${\mathcal{O}}(1)$, and $f$ is a some nonlinear function. In the following, we denote $\mathbf{\boldsymbol{\theta}}=(\mathbf{\boldsymbol{w}},B)$ for the shorthand notation of the linear model’s parameter.
To obtain a classifier, we consider minimizing the following randomly reweighted empirical risk function:
$\hat{\mathbf{\boldsymbol{\theta}}}(\{c_{\mu}^{+}\},\{c_{\nu}^{-}\},D)=\operatorname*{arg\,min}_{(\mathbf{\boldsymbol{w}},b)\in\mathbb{R}^{N+1}}$ $\sum_{\mu=1}^{M^{+}}c_{\mu}^{+}l^{+}(\mathbf{\boldsymbol{x}}_{\mu}^{+}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+b)$
$+\sum_{\nu=1}^{M^{-}}c_{\nu}^{-}l^{-}(\mathbf{\boldsymbol{x}}_{\mu}^{-}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+b)+\lambda\sum_{i=1}^{N}r(w_{i}),$ (4)
when the bias $B$ is estimated by minimizing the cost function, and
$\hat{\mathbf{\boldsymbol{w}}}(\{c_{\mu}^{+}\},\{c_{\nu}^{-}\},D)=\operatorname*{arg\,min}_{\mathbf{\boldsymbol{w}}\in\mathbb{R}^{N}}$ $\sum_{\mu=1}^{M^{+}}c_{\mu}^{+}l^{+}(\mathbf{\boldsymbol{x}}_{\mu}^{+}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+\bar{B})$
$+\sum_{\nu=1}^{M^{-}}c_{\nu}^{-}l^{-}(\mathbf{\boldsymbol{x}}_{\mu}^{-}\cdot\mathbf{\boldsymbol{w}}/\sqrt{N}+\bar{B})+\lambda\sum_{i=1}^{N}r(w_{i}),$ (5)
when the bias is fixed as $\bar{B}$ during the training. Here, $l^{+}$ and $l^{-}$ are the convex loss for positive and negative examples, respectively. For example, the choice of $l^{+}(x)=-\log(\sigma(x))$ and $l^{-}(x)=-\log(1-\sigma(x))$, with $\sigma(\cdot)$ being the sigmoid function, yields the cross-entropy loss. Also, $c_{\mu}^{+}\sim_{\rm iid}p_{c}^{+}$ and $c_{\nu}^{-}\sim_{\rm iid}p_{c}^{-}$ are the random coefficients that represent the effect of the resampling or the weighting. In the following we use the notation $\mathbf{\boldsymbol{c}}\equiv\{c_{\mu}^{+}\}_{\mu=1}^{M^{+}}\cup\{c_{\nu}^{-}\}_{\nu=1}^{M^{-}}$. Finally, $\lambda r(\cdot)$ is the convex regularizer for the weight vector with a regularization parameter $\lambda\in(0,\infty)$. By appropriately specifying the distribution $p_{c}^{\pm}$, the estimators (4) and (5) can represent estimators for different types of costs for resampled or reweighted data as follows:
  • US without replacement (subsampling):
    $p_{c}^{+}(c^{+})$ $=\delta_{\rm d}(c^{+}-1)$ (6)
    $p_{c}^{+}(c^{-})$ $=\mu_{\ast}\delta_{\rm d}(c^{-}-1)+(1-\mu_{\ast})\delta_{\rm d}(c^{-}),$ (7)
    where $\mu_{\ast}\in(0,1)$ be the resampling rate. $\mu_{\ast}=M^{+}/M^{-}$ is the natural choice to balance the size of the sizes of positive and negative examples.
  • US with replacement (bootstrap):
    $p_{c}^{+}(c^{-})$ $=\delta_{\rm d}(c^{+}-1),$ (8)
    $p_{c}^{-}$ $={\rm Poisson}(c^{-};\mu_{\ast}),$ (9)
    where $\mu_{\ast}\in(0,1)$ be the resampling rate.
  • Simple weighting:
    $p_{c}^{+}(c^{+})$ $=\delta_{\rm d}(c^{+}-\gamma^{+}),$ (10)
    $p_{c}^{-}(c^{-})$ $=\delta_{\rm d}(c^{-}-\gamma^{-}),$ (11)
    where $\gamma^{+},\gamma^{-}\in(0,\infty)$ is the reweighting coefficient for the positive and negative classes. $\gamma^{+}=(M^{+}+M^{-})/(2M^{+}),\gamma^{-}=(M^{+}+M^{-})/(2M^{-})$ is the standard choice used in major ML packages such as the scikit-learn (Pedregosa et al., 2011; King & Zeng, 2001).
The ensemble average of the prediction for the input $\mathbf{\boldsymbol{x}}$ is thus obtained as follows:
$\begin{cases}\mathbb{E}_{\mathbf{\boldsymbol{\mathbf{\boldsymbol{c}}}}}\left[\mathbf{\boldsymbol{x}}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}},D)\right],&{\rm when~{}bias~{}is~{}estimated\vspace{2truemm}}\\ \mathbb{E}_{\mathbf{\boldsymbol{\mathbf{\boldsymbol{c}}}}}\left[\mathbf{\boldsymbol{x}}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\bar{B}\right],&{\rm when~{}bias~{}is~{}fixed}\end{cases}$ (12)
In this study, we focus on the case of subsampling as a method of resampling, since the the computational cost is usually smaller than the bootstrap.
The aim of this study is twofold: one is to characterize how the estimators (4) and (5) statistically depend on $D$ and $\mathbf{\boldsymbol{c}}$, and the other is to clarify the difference in performance with the UB obtained by (12) and the US or SW obtained by (4) or (5) with fixed $\mathbf{\boldsymbol{c}}$. We are interested in the generalization performance for both positive and negative inputs. Therefore, we will evaluate the generalization performance using the $F$-measure, the weighted harmonic mean of specificity and recall.
To characterize the properties of the estimators (4) and (5), we consider the large system limit where $N,M^{+},M^{-}\to\infty$, keeping their ratios as $(M^{+}/N,M^{-}/N)=(\alpha^{+},\alpha^{-})\in(0,\infty)^{2}$. In this asymptotic limit, the behavior of the estimators (4) and (5) can be sharply characterized as shown in the next section. We term this asymptotic limit as large system limit (LSL). Hereafter, $N\to\infty$ represents LSL as a shorthand notation to avoid cumbersome notation.
3 Asymptotics of the estimator for the randomized cost function
In this section, we present the asymptotic properties of the estimators (4) and (5) for the randomized cost functions. It is described by a small finite set of scalar quantities determined as a solution of nonlinear equations, which we refer to as self-consistent equations.
First, let us define the self-consistent equations that determines a set of scalar quantities $B\in\mathbb{R},\Theta=(q,m,\chi,v)\in(0,\infty)\times\mathbb{R}\times(0,\infty)\times(0,\infty)$ and $\hat{\Theta}=(\hat{Q},\hat{m},\hat{\chi},\hat{v})\in(0,\infty)\times\mathbb{R}\times(0,\infty)\times(0,\infty)$ that characterize the statistical behavior of (4) and (5).
Claim 1 (Self-consistent equations)
The set of quantities $\Theta$ and $\hat{\Theta}$ are given as the solution of the following set of non-linear equations that are referred to as self-consistent equations. Let $\hat{\mathsf{w}}$ and $\hat{\mathsf{u}}^{\pm}$ be the solution of the following one-dimensional randomized optimization problems:
$\hat{\mathsf{w}}$ $=\operatorname*{arg\,min}_{w\in\mathbb{R}}\frac{\hat{Q}}{2}w^{2}-h_{w}w+\lambda r(w),$ (13)
$\hat{\mathsf{u}}^{\pm}$ $=\operatorname*{arg\,min}_{u\in\mathbb{R}}\frac{u^{2}}{2\chi\Delta}+c^{\pm}l^{\pm}(u+h_{u}^{\pm}),$ (14)
where
$h_{w}$ $=\hat{m}+\sqrt{\hat{\chi}}\xi_{w}+\sqrt{\hat{v}}\eta_{w},\quad\xi_{w},\eta_{w}\sim{\mathcal{N}}(0,1),$ (15)
$h_{u}^{\pm}$ $=B\pm m+\sqrt{q}\xi_{u}+\sqrt{v}\eta_{u},\quad\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta),$ (16)
and $c^{+}\sim p_{c}^{+},c^{-}\sim p_{c}^{-}$. Then, the self-consistent equations are given as follows:
$\chi$ $=E_{\xi_{w},\eta_{w}\sim{\mathcal{N}}(0,1)}\left[\frac{d}{dh_{w}}\hat{\mathsf{w}}\right],$ (17)
$q$ $=\mathbb{E}_{\xi_{w}\sim{\mathcal{N}}(0,1)}\left[\mathbb{E}_{\eta_{w}\sim{\mathcal{N}}(0,1)}[\hat{\mathsf{w}}]^{2}\right],$ (18)
$m$ $=E_{\xi_{w},\eta_{w}\sim{\mathcal{N}}(0,1)}\left[\hat{\mathsf{w}}\right],$ (19)
$v$ $=\mathbb{E}_{\xi_{w}\sim{\mathcal{N}}(0,1)}\left[{\mathbb{V}}_{\eta_{w}\sim{\mathcal{N}}(0,1)}[\hat{\mathsf{w}}]\right],$ (20)
$\hat{Q}$ $=-\frac{1}{\chi}\mathbb{E}_{\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{\pm}\sim p_{c}^{\pm}}\left[\alpha^{+}\frac{d}{dh_{u}^{+}}\hat{\mathsf{u}}^{+}+\alpha^{-}\frac{d}{dh_{u}^{-}}\hat{\mathsf{u}}^{-}\right],$ (21)
$\hat{\chi}$ $=\frac{1}{\Delta\chi^{2}}\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\alpha^{+}\mathbb{E}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{+}\sim p_{c}^{+}}[\hat{\mathsf{u}}^{+}]^{2}+\alpha^{-}\mathbb{E}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{-}\sim p_{c}^{-}}[\hat{\mathsf{u}}^{-}]^{2}\right],$ (22)
$\hat{m}$ $=\frac{1}{\Delta\chi}\mathbb{E}_{\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{\pm}\sim p_{c}^{\pm}}\left[\alpha^{+}\hat{\mathsf{u}}^{+}-\alpha^{-}\hat{\mathsf{u}}^{-}\right],$ (23)
$\hat{v}$ $=\frac{1}{\Delta\chi^{2}}\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\alpha^{+}{\mathbb{V}}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{+}\sim p_{c}^{+}}[\hat{\mathsf{u}}^{+}]+\alpha^{-}{\mathbb{V}}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{-}\sim p_{c}^{-}}[\hat{\mathsf{u}}^{-}]\right].$ (24)
Furthermore, depending on whether the bias is estimated or fixed, the following equation is added to the self-consistent equation
$\begin{cases}0=\mathbb{E}_{\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta),c^{\pm}\sim p_{c}^{\pm}}\left[\alpha^{+}\hat{\mathsf{u}}^{+}+\alpha^{-}\hat{\mathsf{u}}^{-}\right],&{\rm when~{}bias~{}is~{}estimated\vspace{2truemm}}\\ B=\bar{B},&{\rm when~{}bias~{}is~{}fixed}\end{cases}.$ (25)
It is also noteworthy to comment on the interpretation of the parameter $B$ and $(q,m,v)$ in the self-consistent equations.
Claim 2 (Interpretation of the parameters $\Theta$)
At the LSL, the solution of the self-consistent equations are related with the weight vector $\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)$ as follows:
$q$ $=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{\boldsymbol{c}}}\left[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)\right]^{2},$ (26)
$m$ $=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{\mathbf{\boldsymbol{c}}}\left[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)\right],$ (27)
$v$ $=\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}{\mathbb{V}}_{\mathbf{\boldsymbol{c}}}\left[\hat{w}_{i}(\mathbf{\boldsymbol{c}},D)\right],$ (28)
$B$ =$=\begin{cases}\lim_{N\to\infty}\hat{B}(\mathbf{\boldsymbol{c}},D),&{\rm when~{}bias~{}is~{}estiamted\vspace{2truemm}}\\ \bar{B},&{\rm when~{}bias~{}is~{}fixed}\end{cases}$ (29)
That is, $q,m,v,B$ represents the squared norm of the averaged weight vector, the inner product with the cluster center, the variance with respect to the resampling, and the bias, respectively.
Using the solution of the self-consistent equations, we can obtain the following expression for the average of the prediction.
Claim 3 ($l$-th moments of the predictor)
Let $(B,m,q,v)$ be the solution of the self-consistent equation in Claim 1. Also, let $l=0,1,\dots\in{\mathbb{N}}$ be the non-negative integer. Also, $\hat{u}^{\pm}(\mathbf{\boldsymbol{c}},D)\equiv\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\hat{B}(\mathbf{\boldsymbol{c}},D)$ and $\hat{u}^{\pm}(\mathbf{\boldsymbol{c}},D)\equiv\mathbf{\boldsymbol{x}}^{\pm}\cdot\hat{\mathbf{\boldsymbol{w}}}(\mathbf{\boldsymbol{c}},D)/\sqrt{N}+\bar{B}$ be the predictions of the estimators (4) and (5) for an input $\mathbf{\boldsymbol{x}}^{\pm}$ which is generated as
$\mathbf{\boldsymbol{x}}^{\pm}=\pm\mathbf{\boldsymbol{1}}_{N}/\sqrt{N}+\mathbf{\boldsymbol{z}},\quad z_{i}\sim_{\rm iid}p_{z},i\in[N].$ (30)
Then, for any $l$, we can obtain the following result regardless of whether the bias is estimated or not:
$\mathbb{E}_{D,\mathbf{\boldsymbol{x}}^{\pm}}\left[\mathbb{E}_{\mathbf{\boldsymbol{c}}}[\psi(\hat{u}^{\pm})]^{l}\right]=\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\mathbb{E}_{\eta_{u}\sim{\mathcal{N}}(0,\Delta)}[\psi(B\pm m+\sqrt{q}\xi_{u}+\sqrt{v}\eta_{u})]^{l}\right],$ (31)
where $\psi$ is arbitrary as long as the above integrals are convergent.
This result indicates that the fluctuation with respect to resampling conditioned by $(\mathbf{\boldsymbol{x}},y)$ is effectively described by the Gaussian random variable $\eta_{u}$, and the fluctuation with respect to the noise term in $\mathbf{\boldsymbol{x}}$ is described by the Gaussian random variable $\xi_{u}$. Schematically, this behavior can be represented as follows:
$\hat{u}^{\pm}=\underset{\rm bias}{\uwave{B}}+\underset{\begin{subarray}{c}\rm correlation~{}with\\ \rm the~{}cluster~{}center\end{subarray}}{\uwave{(\pm m)}}+\underset{\begin{subarray}{c}{\rm fluctuation~{}from}\\ {\rm the~{}noise~{}}\mathbf{\boldsymbol{z}}\end{subarray}}{\uwave{\sqrt{q}\xi_{u}}}+\underset{\begin{subarray}{c}{\rm fluctuation~{}from}\\ {\rm random~{}reweighting~{}}\mathbf{\boldsymbol{c}}\\ {\rm conditioned~{}by~{}}\mathbf{\boldsymbol{x}}\end{subarray}}{\uwave{\sqrt{v}\eta_{u}}}.$ (32)
Therefore, the distribution of the prediction averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$ can be described as follows.
Claim 4 (Distribution of the averaged prediction)
Let $(B,m,q,v)$ be the solution of the self-consistent equation in Claim 1. Then, the distribution of the prediction $\hat{u}_{K}^{\pm}$ averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$, when the noise term $\mathbf{\boldsymbol{z}}$ fluctuates, is equal to that of the following random quantity:
$\hat{u}^{\pm}_{K}=B\pm m+\sqrt{q}\xi_{u}+\sqrt{\frac{v}{K}}\eta_{u},\quad\xi_{u},\eta_{u}\sim{\mathcal{N}}(0,\Delta).$ (33)
(a) $(\frac{M^{+}}{M},y,N)=(0.01,-1,2^{10})$
(b) $(\frac{M^{+}}{M},y,N)=(0.05,-1,2^{10})$
(c) $(\frac{M^{+}}{M},y,N)=(0.2,-1,2^{10})$
(d) $(\frac{M^{+}}{M},y,N)=(0.01,-1,2^{13})$
(e) $(\frac{M^{+}}{M},y,N)=(0.05,-1,2^{13})$
(f) $(\frac{M^{+}}{M},y,N)=(0.2,-1,2^{13})$
(g) $(\frac{M^{+}}{M},y,N)=(0.01,+1,2^{10})$
(h) $(\frac{M^{+}}{M},y,N)=(0.05,+1,2^{10})$
(i) $(\frac{M^{+}}{M},y,N)=(0.2,+1,2^{10})$
(j) $(\frac{M^{+}}{M},y,N)=(0.01,+1,2^{13})$
(k) $(\frac{M^{+}}{M},y,N)=(0.05,+1,2^{13})$
(l) $(\frac{M^{+}}{M},y,N)=(0.2,+1,2^{13})$
Figure 1: Comparison between the empirical distribution of $\hat{u}_{K}^{\pm}$ (histogram), which is obtained by a single realization of the training data $D$ of finite size, and the theoretical prediction in Claim 4 (dashed). Different colors represent resampling averages with different numbers of realizations of reweighting coefficient $\mathbf{\boldsymbol{c}}$.
Using the above result, one obtains the expression of the $F$-measure as follows:
Claim 5 ($F$-measure)
Let $(m,q,B)$ be the solution of the self-consistent equations in Claim 1. Then, for the prediction $\hat{u}_{K}^{\pm}$ averaged over $K$-realization of the reweighing coefficient $\{\mathbf{\boldsymbol{c}}_{k}\}_{k=1}^{K}$, specificity ${\mathcal{S}}$, recall ${\mathcal{R}}$, and the $F$-measure ${\mathcal{F}}$ that is obtained as the weighted harmonic mean of specificity and recall, is given as follows:
${\mathcal{S}}$ $=H\left(-\frac{m+B}{\sqrt{\Delta(q+\frac{v}{K})}}\right),$ (34)
${\mathcal{R}}$ $=1-H\left(-\frac{-m+B}{\sqrt{\Delta(q+\frac{v}{K})}}\right),$ (35)
${\mathcal{F}}$ $=\frac{2}{\frac{1}{{\mathcal{S}}}+\frac{1}{{\mathcal{R}}}},$ (36)
where $H(\cdot)$ is the complementary error function:
$H(x)=\mathbb{E}_{\xi\sim{\mathcal{N}}(0,1)}\left[\mbox{\rm 1}\mbox{\rm l}(\xi>x)\right].$ (37)
If the bias is properly estimated, the generalization performance is determined by the variance of the prediction $\Delta(q+v/K)$.
3.1 Cross-checking with numerical experiments
To check the validity of Claim 4, which is the most important result for characterizing the generalization performance, we briefly compare the result of numerical solution of the self-consistent equations with numerical experiments of finite-size systems.
We consider the case of subsampling (6)-(7). For the loss function, we use the cross entropy loss; $l^{+}(x)=-\log(\sigma(x))$ and $l^{-}(x)=-\log(1-\sigma(x))$, where $\sigma(\cdot)$ is the sigmoid function. The regularizer is the ridge: $r(x)=x^{2}/2$. Also, the bias is estimated by minimizing the loss as in (4). In experiments, the scikit-learn package is used (Pedregosa et al., 2011), and the system sizes $N=2^{10}$ and $N=2^{13}$ are used. For simplicity, we fix the regularization parameter as $\lambda=10^{-4}$, the size of the training data as $(M^{+}+M^{-})/N=1/2$, and the variance of the noise as $\Delta=0.75^{2}$. For evaluating the distribution of $\hat{u}_{K}^{\pm}$ empirically, we use $10^{5}$ examples of $\mathbf{\boldsymbol{x}}^{\pm}$.
Figure 1 shows the comparison the distribution in Claim 4 and the empirical distribution obtained by experiments with finite $N$. Different colors corresponds to different numbers of samples for the resampling average. Each panel corresponds to a different set of $(M^{+}/M,y,N)$, where $y$ is the label to indicate positive ($y=1$) or negative ($y=-1$) classes. In all cases, although some discrepancies between theoretical prediction and experiment are observed when $N=2^{10}$, they are in agreement when $N=2^{13}$, demonstrating the validity of Claim 4.
4 Under-bagging and regularization
(a) $(\sqrt{\Delta},\lambda)=(1.5,10^{-3})$
(b) $(\sqrt{\Delta},\lambda)=(1.5,10^{-1})$
(c) $(\sqrt{\Delta},\lambda)=(1.5,10^{0})$
(d) $(\sqrt{\Delta},\lambda)=(0.75,10^{-3})$
(e) $(\sqrt{\Delta},\lambda)=(0.75,10^{-1})$
(f) $(\sqrt{\Delta},\lambda)=(0.75,10^{0})$
(g) $(\sqrt{\Delta},\lambda)=(0.5,10^{-3})$
(h) $(\sqrt{\Delta},\lambda)=(0.5,10^{-1})$
(i) $(\sqrt{\Delta},\lambda)=(0.5,10^{0})$
Figure 2: The comparison of the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and the minority (positive) classes, between US (dashed) and UB (solid). Each panel corresponds to different values of $(\Delta,\lambda)$. Different colors of the lines represent the result for different sizes of the minority classes $\alpha^{+}$.
In this section, using Claim 4, we compare the performance of several estimators, that is, (i) US, (ii) UB, and (iii) SW.
For simplicity, we focus on the case of ridge-regularized cross-entropy loss as in subsection 3.1. When training a model using an under-sampled data set as in (6)-(7), we can expect that the estimated bias term $\hat{B}(\mathbf{\boldsymbol{c}},D)$ to be zero, since the number of positive and negative samples is balanced. Therefore, we set the bias as zero a priori when using subsampled data sets. Furthermore, as investigated in the previous studies (Dobriban & Wager, 2018; Mignacco et al., 2020), when training a linear classifier using balanced data and a ridge-regularized cross-entropy loss, an infinitely large large regularization parameter yields the best direction for the classification plane. However, as reported in (Mignacco et al., 2020; Takahashi, 2024), the discrepancy between the theory and experiment with finite N can be huge if too a large regularization parameter is used. Therefore, for US and UB, we consider several finite regularization parameters for US and UB. When considering the weighting, the behavior of bias and the optimal regularization parameter, the bias term is estimated as in (4) and the regularization parameter is optimized to maximize the $F$-measure using the Nelder-Mead method in Optim.jl package (Mogensen & Riseth, 2018).
4.1 US and UB
In this subsection, we compare the performance of US and UB, through the values of $F$-measures.
Figure 2 shows the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and minority (positive) classes. Each panel corresponds to different values of the variance $\Delta$ and the regularization parameter $\lambda$. Recall that the $F$-measure is uppder bounded by unity. In UB, the $F$-measure increases monotonically as the number of excess data points in the majority class $\alpha^{-}-\alpha^{+}$ increases, despite of the large class imbalance. However, in the case of US, the $F$-measure does not depends on $\alpha^{-}-\alpha^{+}$, i.e., it is not affected by the class imbalance, but cannot utilize the excess data points in the majority class. This tendency does not depend on the values of $(\Delta,\lambda)$. Actually, when subsampling is used as the resampling method, one can show that the performance of the classifier does not depends on the $\alpha^{-}-\alpha^{+}$ as shown below.
The $F$-measure for the estimator obtained by US, which corresponds to the case of $K=1$ in Claim 5, is determined by $\bar{q}=q+v$. Let us also define $\hat{\bar{\chi}}=\hat{\chi}+\hat{v}$. Then, by adding both side of equations (18) and (20) as well as (22) and (24), and taking the average of $c^{\pm}\sim p_{c}^{\pm}$ explicitly, the following modified self-consistent equations that determine $\bar{q},m,B$ are obtained as follows.
Claim 6 (Modified self-consistent equations for US)
Let $\bar{q}=q+v$ and $\hat{\bar{\chi}}=\hat{\chi}+\hat{v}$. Then, the set of quantities $\bar{q},m,\chi$ and $\hat{Q},\hat{\bar{\chi}},\hat{m}$ are given as the solution of the following set of modified self-consistent equations. Let $\bar{\mathsf{w}}$ and $\bar{\mathsf{u}}^{\pm}$ be the solution of the following one-dimensional randomized optimization problems:
$\bar{\mathsf{w}}$ $=\operatorname*{arg\,min}_{w\in\mathbb{R}}\frac{\hat{Q}}{2}w^{2}-\bar{h}_{w}w+\lambda r(w),$ (38)
$\bar{\mathsf{u}}^{\pm}$ $=\operatorname*{arg\,min}_{u\in\mathbb{R}}\frac{u^{2}}{2\chi\Delta}+l^{\pm}(u+\bar{h}_{u}^{\pm}),$ (39)
where
$\bar{h}_{w}$ $=\hat{m}+\sqrt{\hat{\chi}}\xi_{w},\quad\xi_{w},\sim{\mathcal{N}}(0,1),$ (40)
$\bar{h}_{u}^{\pm}$ $=B\pm m+\sqrt{\bar{q}}\xi_{u},\quad\xi_{u}\sim{\mathcal{N}}(0,\Delta).$ (41)
Then, the modified self-consistent equations are given as follows:
$\chi$ $=E_{\xi_{w},\sim{\mathcal{N}}(0,1)}\left[\frac{d}{d\bar{h}_{w}}\bar{\mathsf{w}}\right],$ (42)
$\bar{q}$ $=\mathbb{E}_{\xi_{w}\sim{\mathcal{N}}(0,1)}\left[\bar{\mathsf{w}}^{2}\right],$ (43)
$m$ $=E_{\xi_{w}\sim{\mathcal{N}}(0,1)}\left[\bar{\mathsf{w}}\right],$ (44)
$\hat{Q}$ $=-\frac{\alpha^{+}}{\chi}\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\frac{d}{d\bar{h}_{u}^{+}}\bar{\mathsf{u}}^{+}+\frac{d}{d\bar{h}_{u}^{-}}\bar{\mathsf{u}}^{-}\right],$ (45)
$\hat{\bar{\chi}}$ $=\frac{\alpha^{+}}{\Delta\chi^{2}}\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[(\bar{\mathsf{u}}^{+})^{2}+(\bar{\mathsf{u}}^{-})^{2}\right],$ (46)
$\hat{m}$ $=\frac{\alpha^{+}}{\Delta\chi}\mathbb{E}_{\xi_{u},\sim{\mathcal{N}}(0,\Delta)}\left[\bar{\mathsf{u}}^{+}-\bar{\mathsf{u}}^{-}\right].$ (47)
Furthermore, depending on whether the bias is estimated or fixed, the following equation is added to the modified self-consistent equation
$\begin{cases}0=\mathbb{E}_{\xi_{u}\sim{\mathcal{N}}(0,\Delta)}\left[\bar{\mathsf{u}}^{+}+\bar{\mathsf{u}}^{-}\right],&{\rm when~{}bias~{}is~{}estimated\vspace{2truemm}}\\ B=\bar{B},&{\rm when~{}bias~{}is~{}fixed}\end{cases}.$ (48)
It is clear that the above modified self-consistent equation only depends on $\alpha^{+}$ and does not depends on $\alpha^{-}-\alpha^{+}$. Therefore, the $F$-measure of US cannot be improved by the excess examples the majority class111 This is exactly the same with the generalization error for the linear classifier for the data generated from symmetric Gaussian mixture model obtained in (Mignacco et al., 2020). .
(a) $(\sqrt{\Delta},\lambda)=(1.5,10^{-5})$
(b) $(\sqrt{\Delta},\lambda)=(1.5,10^{-1})$
(c) $(\sqrt{\Delta},\lambda)=(1.5,10^{0})$
(d) $(\sqrt{\Delta},\lambda)=(0.75,10^{-5})$
(e) $(\sqrt{\Delta},\lambda)=(0.75,10^{-1})$
(f) $(\sqrt{\Delta},\lambda)=(0.75,10^{0})$
(g) $(\sqrt{\Delta},\lambda)=(0.5,10^{-5})$
(h) $(\sqrt{\Delta},\lambda)=(0.5,10^{-1})$
(i) $(\sqrt{\Delta},\lambda)=(0.5,10^{0})$
Figure 3: The heatmap plot for the relative log-$F$-measure $\log_{10}({\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm US})$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm US}$ are the $F$-measure for UB and US, respectively. Each panel corresponds to different values of $(\Delta,\lambda)$.
Figure 3 shows heatmap of the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm US}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm US}$ are the $F$-measure for UB and US, respectively. It is clear that the values of the $F$-measures are improved when the variance is small, the size of the minority class is small and the excess size of the majority class is large. When the variance is large, the improvement at small $\alpha^{+}$ region is small because $m$ in (36) can be small due to the large overlap between two clusters. It is also noteworthy that the improvement may be large around some values of $\alpha^{+}$ when the regularization parameter is extremely small. This indicates that the UB is robust to the existence of the phase transition from linearly separable to inseparable training data, although the performance of standard interpolator obtained by a single realization of training data is affected by such a phase transition as reported in (Mignacco et al., 2020).
4.2 UB and SW
(a) $(\sqrt{\Delta},\lambda)=(1.5,10^{-5})$
(b) $(\sqrt{\Delta},\lambda)=(1.5,10^{-1})$
(c) $(\sqrt{\Delta},\lambda)=(1.5,10^{0})$
(d) $(\sqrt{\Delta},\lambda)=(0.75,10^{-5})$
(e) $(\sqrt{\Delta},\lambda)=(0.75,10^{-1})$
(f) $(\sqrt{\Delta},\lambda)=(0.75,10^{0})$
(g) $(\sqrt{\Delta},\lambda)=(0.5,10^{-5})$
(h) $(\sqrt{\Delta},\lambda)=(0.5,10^{-1})$
(i) $(\sqrt{\Delta},\lambda)=(0.5,10^{0})$
Figure 4: The comparison of the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and the minority (positive) classes, between weighting (dashed) and UB (solid). Each panel corresponds to different values of $(\Delta,\lambda)$. Different colors of the lines represent the result for different sizes of the minority classes $\alpha^{+}$.
In this subsection, we compare the performance of UB with SW, through the values of $F$-measures. For SW, the standard weighting coefficients $\gamma^{+}=(\alpha^{+}+\alpha^{-})/(2\alpha^{+}),\gamma^{-}=(\alpha^{+}+\alpha^{-})/(2\alpha^{-})$ are used.
Figure 4 shows the dependence of the $F$-measure ${\mathcal{F}}$ on $\alpha^{-}-\alpha^{+}$, which is the difference in the size between the majority (negative) and minority (positive) classes. In contrast to US and UB, the $F$-measure of the SW method degrades as the number of the excess examples of the majority class increases. In other words, as the amount of the class imbalance increases, the prediction performance of the estimator obtained by simple class-wise weighted cost degrades, although the degradation is mild if the size of the minority class is large. Similarly, Figure 4 shows heatmap of the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm SW}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm SW}$ are the $F$-measure for UB and SW, respectively. When the size of the minority class is sufficiently large, the performance is almost comparable. However, when the size of the minority class is small and the class imbalance is large, the performance of the simple weighting method is several times to tens of thousands of times worser than UB in terms of the value of $F$-measure.
(a) $(\sqrt{\Delta},\lambda)=(1.5,10^{-5})$
(b) $(\sqrt{\Delta},\lambda)=(1.5,10^{-1})$
(c) $(\sqrt{\Delta},\lambda)=(1.5,10^{0})$
(d) $(\sqrt{\Delta},\lambda)=(0.75,10^{-5})$
(e) $(\sqrt{\Delta},\lambda)=(0.75,10^{-1})$
(f) $(\sqrt{\Delta},\lambda)=(0.75,10^{0})$
(g) $(\sqrt{\Delta},\lambda)=(0.5,10^{-5})$
(h) $(\sqrt{\Delta},\lambda)=(0.5,10^{-1})$
(i) $(\sqrt{\Delta},\lambda)=(0.5,10^{0})$
Figure 5: The heatmap plot for the relative $F$-measure ${\mathcal{F}}_{\rm UB}/{\mathcal{F}}_{\rm weighting}$, where ${\mathcal{F}}_{\rm UB}$ and ${\mathcal{F}}_{\rm weighting}$ are the $F$-measure for UB and the weighting method, respectively. Each panel corresponds to different values of $(\Delta,\lambda)$.
5 Summary and Discussions
In this work, we derived a sharp asymptotics of the estimators obtained by the randomly reweighted losses (4) and (5) in the limit where the data size and the input dimension diverge proportionally. The derivation was based on the standard use of the replica method. The results are summarized in Claim 1-5.
By using the derived sharp asymptotics, we investigated the performance of the UB, US and SW. Our main findings were (i) UB can improve the performance of the classifier in terms of the $F$-measure even if the class-imbalance is huge, (ii) the performance of the classifier obtained by US does not depends on the size of the excess majority examples, and (iii) the performance of the SW method degrades as the size of the excess majority examples increases especially when the size of the minority class is small and the class imbalance is large. Furthermore, UB seemed to be robust to the interpolation phase transition in contrast to US.
Although UB is a powerful method for learning in imbalanced data, but it requires an increased computational cost proportional to the number of under-sampled datasets, which can be a serious issue in learning deep neural networks. Therefore, developing simple heuristics to achieve similar results as UB would be a promising future direction.
Acknowledgments
This study was supported by JSPS KAKENHI Grant No. 21K21310 and 23K16960, and Grant-in-Aid for Transformative Research Areas (A), “Foundation of Machine Learning Physics” (22H05117).